A stroboscope is flashing light signals at the rate of 2100 flashes/min. An observer in a rocket ship traveling toward the strobe light at 0.500c would see what flash rate?
1
Expert's answer
2020-09-28T06:06:12-0400

speed of Rocket=0.500c

Rate of light signal=2100 flahes/min


Using time dilation equation

"t=\frac{t_o}{\sqrt{1-\frac{v^2}{c^2}}}"

="\frac{1}{\sqrt{1-\frac{0.25c^2}{c^2}}}"


="\frac{1}{\sqrt{0.75}}"


=1.154 min


Number .of flashes seen in 1.154 min="1.154\times 2100"

=2423.4

i.e. approx 2432 flashes/min